# Mechanically Programming the Cross-Sectional Shape of Soft Growing Robotic Structures for Patient Transfer

### v1

@@ Abstract
Pneumatic soft everting robotic structures have the potential to facilitate human transfer tasks due to their ability to grow underneath humans without sliding friction and their utility as a flexible sling when deflated. Tubular structures naturally yield circular cross-sections when inflated, whereas a robotic sling must be both thin enough to grow between them and their resting surface and wide enough to cradle the human. Recent works have achieved flattened cross-sections by including rigid components into the structure, but this reduces conformability to the human. We present a method of mechanically programming the cross-section of soft everting robotic structures using flexible strips that constrain radial expansion between points along the outer membrane. Our method enables simultaneously wide and thin profiles while maintaining the full multi-axis flexibility of traditional slings. We develop and validate a model relating the geometric design specifications to the fabrication parameters, and experimentally characterize their effects on growth rate. Finally, we prototype a soft growing robotic sling system and demonstrate its use for assisting a single caregiver in bed-to-chair patient transfer.
@@ I INTRODUCTION
The field of soft robotics continues to realize the promise of leveraging mechanically compliant robotic designs as technological solutions to emerging societal challenges [1]. One potential application of such work lies in securely harnessing and transferring humans. In eldercare and the care for people with physical disabilities, transferring humans is a critical task regularly carried out by caregivers [2]. This removes the patient’s independence in daily life and can be strenuous and cause injuries for caregivers [3, 4, 5]. The current standard practice for caregivers is to manually place the patient in a sling or harness to securely lift them [2]. Slings and straps are practical for high-force human interaction because they can be wrapped around the body to bear the heavy load while distributing it over a large contact area [6]. Slings are particularly beneficial, in that their wide sheet-like topology enables maximal contact area with the body. In current practice, caregivers must manually place the patient into and out of the sling, inflicting harmful strain on their body [3]. Such manual patient handling operations are fatiguing, dangerous, and require multiple caregivers.
@@
While previous robots have been developed for patient transfer tasks [7, 8, 9], traditional robotic manipulators cannot provide the simultaneously strong yet gentle interaction necessary to safely lift the full weight of the human body. Inspired by straps and slings used in eldercare, a semi-soft hyper-redundant robot with the high tensile strength and bending flexibility needed to gently bear high loads was previously developed, and demonstrated lifting the full weight of humans without external physical assistance [6]. However, robotic manipulators with any rigid components cannot easily be placed between the patient and their resting surface without an external party adjusting and/or lifting their body, primarily due to the lack of a gap to enter and the resulting sliding friction. Recent literature has demonstrated soft robotic designs leveraging eversion to grow (and retract from) under the patient without sliding friction [10, 11]. [10] utilizes pneumatically driven eversion to insert soft cylindrical manipulators underneath the human body to aid with turning. This work is limited to a partial assist as opposed to a full body patient transfer. Towards realizing the employing of soft everting structures for full body transfer, [11] presented a pneumatically driven growing sling that restricted the radial expansion of an inflated beam to achieve a flatter shape under the patient, better suited for their comfort, by integrating rigid shafts into the membrane. These rigid shafts restrict the sling’s flexibility to bending only about the patient’s frontal axis, and limit the cross-sectional profile to convex shapes.
@@
In this paper, we present a method for mechanically programming the cross-sectional shape of soft, everting, compliant robotic structures and realize a self-tunneling and retracting sling. We present the first demonstration of safely harnessing and transferring the full weight of a human with a robotic sheet-shaped sling device used by a single operator. Our robotic sling extends and retracts under the body automatically with no sliding friction, eliminating the uncomfortable experience of being rolled back and forth. The design maintains high global all-axis bending compliance with high tensile integrity by employing a simple fabrication process that includes only heat-sealable fabric. We also present and validate a model to select the fabrication parameters for a desired cross-section design. Fabrication methods are developed to realize our design approach, and a prototype of a soft growing pneumatic sling with full multi-axis bending flexibility is designed and built for patient transfers. The system is experimentally characterized, demonstrating its ability to harness and lift humans effectively for bed-to-chair transfers.
@@ II-A1 Eversion
Our system achieves “growth” and “retraction” at its tip through pressure-driven eversion of flexible, thin-walled membranes. This process, first presented in [12], refers to starting with the flexible, thin-walled, inextensible membranes inverted and achieving lengthening from the tip by pressuring the system to transfer material from within its tubular body to the outside. The growth rate of the robotic structure is additionally constrained by a “tail” (e.g. a string) attached to the end of its everted material. Control of the internal pressure and string release is required for precise control of the system’s growth. This eversion principle coupled with the compliance of the membranes enables soft growing robots to achieve high extension ratios and navigate through cluttered environments without external surface friction. Specific to our application, it allows for navigation through gaps that are significantly smaller than its width, which is required to insert a sling underneath a patient in a supine position. By using eversion, the flexible membrane can grow into small gaps between the subject’s body and the bed surface without inducing damaging friction to the skin and wounding the patient. The lifting force needed for soft growing robotic structures leveraging eversion to grow into the slanted gap underneath the patient has been studied in past literature [12, 13, 11], and shown to be primarily dependent on the internal air pressure and the angle of the slanted gap.
@@ II-A2 Restricting Radial Expansion
The ideal geometry for a structure meant to grow underneath patients is one that is wide enough to span the patient’s body and flat enough to avoid discomfort. The inflated geometry of soft everting growing robots, barring external forces, is governed by its wall tension when pressurized. By systematically fabricating the robot body such that it experiences wall tension that restricts its radial expansion into a desired shape, we can design the everting body to have a flatter inflated geometry. The first soft growing robotic structure flat enough to be applicable as a growing sling for patients was described in [11]. This design restricts radial expansion by integrating shafts into the flexible membrane serving as the robot body. The shafts provide an antagonistic moment for the radially expansive forces and are aligned such that they do not interfere with eversion. However, this design results in a loss of all-axis bending flexibility that allows for better conformation of the sling to the body . In our design, we address this by using only flexible fabric. As shown in Fig. 2(a), we bond pairs of points along the robot body’s cross-sectional circumference perpendicular to the direction of its net eversion growth to constrain its radial expansion when inflated. This is akin to how some air mattress designs achieve their flat surface topology. To make this design viable for the eversion principle, the bonded points must allow for a middle channel for the flexible membrane to evert from and sufficiently low capstan friction from the internal materials sliding against each other [13, 14].
@@ II-B1 Fabricating the robot body
To implement restriction of radial expansion, we used thermoplastic-coated fabric (TPU-coated ripstop nylon) due to its exceptional tensile strength and ease of forming robust bonds through heat-sealing the coated sides using an ultrasonic welder (VETRON 5064). The process involved cutting the fabric material that forms the robot body’s perimeter. Subsequently, two strips of fabric were cut to span the net growth length of the robot body, as depicted in Fig. 2(c). Each strip was heat-sealed along two predetermined lines on the perimeter fabric material in the direction of the robot body’s net eversion growth. The distance between these lines was prescribed to constrain the radial expansion by a predetermined amount. The calculation of this distance is detailed in Section III. Finally, the perimeter fabric material was folded to realize the form shown in Fig. 2(c), ensuring that the TPU-coated side contacts itself so it can be heat-sealed along its edges to create an airtight enclosure for the robot body.
@@ II-B2 Creating strong attachments
The fastening points between the Hoyer lift cables and the membrane also creates force concentrations when lifting the human body. While most commercial patient slings have webbing loops sewn onto their edges for the cables to attach to, sewing onto the membrane introduces punctures that lead to air leaks and/or tears. Thus, our prototype implements tubular loops of fabric (same as membrane) welded to the membrane, as shown in Fig. 2(c). The loops are welded inside of the membrane’s outer seam such that the welds are pulled in shear instead of peeling for increased strength.
@@ II-B3 Wide base
The base and collar were designed with a long, narrow shape to match the flattened cross-section of the robotic sling, as well as enable the base to be placed under the human’s head. As shown in Fig. 2(c), a hose clamp fastens a set of curved blocks around the collar, which clamp the robotic sling membrane to the collar. Because the normal pressure the the hose clamp applies at a given point is proportional to its local curvature, the blocks are designed such that the hose clamp holds a roughly elliptical profile (as opposed to the slot-shaped collar) so that it applies some normal pressure around the entire perimeter without excessively protruding from the long and narrow shape of the base.
@@ III-B2 Arcs of center channel are coradial
A governing principle for the geometry of inflated structures is that, assuming inextensibility and no flexural rigidity, the membrane is deformed such that the internal volume is maximized ($A_{c}=A_{c,max}$). Thus, $\frac{dA_{c}}{d\theta_{c}}=0$ must be true for any $S_{c}$, $L$, and $\theta_{c}$ (geometric parameters which fully determine the cross-section geometry). If $2\alpha=L$ must be true for $A_{c}=A_{c,max}$ to be true for any $S_{c}$, $L$, and $\theta_{c}$, then $p_{0,y}=p_{cu,y}=p_{cl,y}=0$. The cross-sectional area and its derivative with respect to $\theta_{c}$ are expressed as:
@@ III-C Cross-Section Geometry Model
Additionally, given the fabrication parameters, the full geometry of the cross section can then be determined. The radii of the curved segments of the side and center channels are given as $r_{s}=H_{s}/2$ and $r_{c}=H_{s}/2$, respectively. The centers of the side arcs are $\mathbf{p_{sr}}=\left[p_{sr,x},p_{sr,y}\right]^{T}$ and $\mathbf{p_{sl}}=\left[p_{sl,x},p_{sl,y}\right]^{T}$, where $p_{sr,x}=\frac{w}{2}-r_{s}$, $p_{sl,x}=-\frac{w}{2}+r_{s}$, and $p_{sr,y}=p_{sl,y}=0$. The centers of the top and bottom arcs are already defined previously as $\mathbf{p_{ct}}=\mathbf{p_{cb}}=\mathbf{p_{0}}=\left[0,0\right]^{T}$. The midpoints of the straight segments are given as: $\mathbf{p_{Lr}}=\left[\frac{w_{c}}{2},0\right]^{T}$ and $\mathbf{p_{Ll}}=\left[-\frac{w_{c}}{2},0\right]^{T}$, where the width of the center channel $w_{c}$ is given as
@@ IV-B Effects of Fabrication Parameters on Growth Rate
Recognizing that this design approach limits the space inside the robot body for new material to evert through thus increasing the contact area between the internal everting material and external everted material and resultantly the capstan friction experienced by the system, we performed this experiment to investigate the effects of the fabrication parameters, $S_{c}$ and $L$ on the growth rate of the soft growing robotic structure at a constant pressure. We fabricated three iterations of the robot body using our design approach holding their perimeter distance at a constant $559$ mm but varying the fabrication parameters holding one of each fabrication parameter constant in at least two of the iterations. The iterations are the same as those used in the geometry model validation experiments. Structure 2 holds the same $S_{c}$ as Structure 1 but a smaller $L$ to study the effects of $L$ on growth rate. Structure 3 holds the same $L$ as Structure 1 but a smaller $S_{c}$ to study the effects of $S_{c}$ on growth rate. To provide a basis of comparison to previously presented soft growing robots without added bonds along their cross-sectional perimeter [12], we also fabricated and tested a tubular body structure of the same perimeter distance as the other structures with an unconstrained radial expansion. Each structure was $914$ mm in length and attached to an airtight pressurized base. We use the same pressure regulator and pressure sensor used in the geometry model validation experiments to control and monitor the robot body internal pressure. As a preliminary step, the robot body is everted to its full length, and then inverted by $432$ mm to establish the starting position. Before each trial, the robot body is inverted back to the starting position. For each trial, the robot growth is physically impeded while it is inflated to a starting pressure of 2.07 kPa. Once the starting pressure is reached, the robot growth is unobstructed and the robot is allowed to evert until it stops. To limit the effects of fabric wrinkling on the growth rate, we chose a pressure level sufficiently high enough for the robot body to consistently grow to complete eversion without stopping. We perform three trials for each robotic structure. Each trial was filmed with a high-definition camera at 30 frames per second, and the video was post-processed using MATLAB’s Computer Vision Toolbox to track the growth rate of the robotic structures. The results are presented in Fig. 4(b).
@@ IV-C System Deployment
To demonstrate the effectiveness of our system, the prototype was used to harness and transfer human subjects from a bed to a chair, and then back to the bed. This demonstration was performed with five participants (5 adult males, age 19-32, weight 56-98 kg). As shown in Fig. 5, with the participant lying prone on a bed, a caregiver placed the base with the robotic sling fully inverted underneath their head. The base was inflated to 34 kPa (5 psi) and the robotic sling was everted to grow underneath the participant. Once grown to the full length of their body, the robotic sling was then deflated to assume a sheet-like form, and the caregiver then attached the loops on the robotic sling to the cables on a Hoyer lift. The Hoyer lift was then operated to lift the patient from the bed and move them into a chair, using standard procedures used in caregiving industry. To reverse the transfer, the participant was lifted from the chair and moved back to lay them onto the bed. The robotic sling was then inflated and inverted to remove it from underneath the participant, after which the caregiver removed the base from under the participant’s head. The study was approved by the Massachusetts Institute of Technology’s Internal Review Board, and participants gave informed consent. The prototype was successful in completing every step of the demonstration for all participants, in that it automatically placed the robotic sling under the participant’s body without human intervention, it enabled the participant to be lifted and transferred to and from a chair with the same procedure used in caregiving industry, and the robotic sling was then automatically removed from under the participant without human intervention.
IV-C1 The Effects of Fabrication Parameters on Patient Comfort
@@
As a method of quantifying the comfort of patients while the robot is deployed underneath them, we define an ergonomic index as the inverse of the slope between the highest point of the $A_{s}$ channel and the highest point of the $A_{c}$ channel (denoted in Fig. 3(a)). A higher ergonomic index trends towards a concave cross-sectional profile of the robotic sling. As intended by the design, the horizontal slot shape allows for a flatter surface for the patient to lay on by minimizing the rounding of the patient’s back that is present with cylindrical soft growing robotic structures. To investigate the effects the fabrication parameters have on the ergonomic index, we use our geometric model to simulate a sweep of ergonomic indices when holding the structure’s perimeter distance constant while varying the fabrication parameters. The results are presented in Fig. 4(c). The data suggests that, for parameter values that yield physically reasonable shapes, decreasing $L_{c}$ while holding $S_{c}$ constant increases the ergonomic index. It also suggests that decreasing $S_{c}$ while holding $L_{c}$ constant increases the ergonomic index. For each combination of $S_{c}$ and $L_{c}$, there exists a cross-sectional perimeter value that maximizes the ergonomic index which practitioners of this design approach can leverage the analytical model to find.

### v3

@@ Abstract
Pneumatic soft everting robotic structures have the potential to facilitate human transfer tasks due to their ability to grow underneath humans without sliding friction and their utility as a flexible sling when deflated. Tubular structures naturally yield circular cross-sections when inflated, whereas a robotic sling must be both thin enough to grow between a human and their resting surface and wide enough to cradle the human. Recent works have achieved flattened cross-sections by including rigid components into the structure, but this reduces conformability to the human. We present a method of mechanically programming the cross-section of soft everting robotic structures using flexible strips that constrain radial expansion between points along the outer membrane. Our method enables simultaneously wide and thin inflated profiles, and maintains the full multi-axis flexibility of traditional slings when deflated. We develop and validate a model relating geometric design specifications to fabrication parameters, and experimentally characterize their effects on growth rate. Finally, we prototype a soft growing robotic sling system and demonstrate its use for assisting a single caregiver in bed-to-chair patient transfer.
@@ I INTRODUCTION
The field of soft robotics continues to realize the promise of leveraging mechanically compliant robotic designs as technological solutions to emerging societal challenges [yasa_overview_2023]. One potential application of such work lies in securely harnessing and transferring humans. In eldercare and the care for people with physical disabilities, transferring humans is a critical task regularly carried out by caregivers [dewit_fundamental_2013]. This removes the patient’s independence in daily life and can be strenuous and cause injuries for caregivers [chang_impact_2010, noauthor_employer-reported_2023, feng_global_2019]. The current standard practice for caregivers is to manually place the patient in a sling or harness to securely lift them [dewit_fundamental_2013]. Slings and straps are practical for high-force human interaction because they can be wrapped around the body to bear the heavy load while distributing it over a large contact area [barhydt_high-strength_2023]. Slings are particularly beneficial, in that their wide sheet-like topology enables maximal contact area with the body. In current practice, caregivers must manually place the patient into and out of the sling, inflicting harmful strain on their body [chang_impact_2010]. Such manual patient handling operations are fatiguing, dangerous, and require multiple caregivers.
@@
While previous robots have been developed for patient transfer tasks [mukai_whole-body_2011, noauthor_tmc_nodate, bolli_enhancing_2024], traditional robotic manipulators cannot provide the simultaneously strong yet gentle interaction necessary to safely lift the full weight of the human body. Inspired by straps and slings used in eldercare, a semi-soft hyper-redundant robot with the high tensile strength and bending flexibility needed to gently bear high loads was previously developed, and demonstrated lifting the full weight of humans without external physical assistance [barhydt_high-strength_2023]. However, robotic manipulators with any rigid components cannot easily be placed between the patient and their resting surface without an external party adjusting and/or lifting their body, primarily due to the lack of a gap to enter and the resulting sliding friction. Recent literature has demonstrated soft robotic designs leveraging eversion to grow (and retract from) under the patient without sliding friction [nakamura_soft_2018, choi_development_2020]. [nakamura_soft_2018] utilizes pneumatically driven eversion to insert soft cylindrical manipulators underneath the human body to aid with turning. This work is limited to a partial assist as opposed to a full body patient transfer. Towards realizing the employing of soft everting structures for full body transfer, [choi_development_2020] presented a pneumatically driven growing sling that restricted the radial expansion of an inflated beam to achieve a flatter shape under the patient, better suited for their comfort, by integrating rigid shafts into the membrane. These rigid shafts restrict the sling’s flexibility to bending only about the patient’s frontal axis, and limit the cross-sectional profile to convex shapes. The authors of this paper have previously demonstrated lifting the full weight of a human with multiple fully soft growing robots without any rigid components [barhydt2025loopclosuregraspingtopological]. However, the growing robots had a cylindrical cross-section and were used as individuals straps working together to support the human along their upper back and under their knees, unlike traditional patient slings which can be wide enough to better securely support the entire body with just a single sheet.
@@
In this paper, we present a method for mechanically programming the cross-sectional shape of soft, everting, compliant robotic structures and realize a self-tunneling and retracting sling. We present the first demonstration of safely harnessing and transferring the full weight of a human with a robotic sheet-shaped sling device used by a single operator. Our robotic sling extends and retracts under the body automatically with no sliding friction, eliminating the uncomfortable experience of being rolled back and forth. Once fully deployed and deflated, the design maintains high global all-axis bending compliance about all body axes (frontal, sagittal, longitudinal) with high tensile integrity by employing a simple fabrication process that includes only heat-sealable fabric. We also present and validate a model to select the fabrication parameters for a desired cross-section design. Fabrication methods are developed to realize our design approach, and a prototype of a soft growing pneumatic sling with full multi-axis bending flexibility is designed and built for patient transfers. The system is experimentally characterized, demonstrating its ability to harness and lift humans effectively for bed-to-chair transfers.
@@ II-A1 Eversion
Our system achieves “growth” and “retraction” at its tip through pressure-driven eversion of flexible, thin-walled membranes. This process, first presented in [hawkes_soft_2017], refers to starting with the flexible, thin-walled, inextensible membranes inverted and achieving lengthening from the tip by pressuring the system to transfer material from within its tubular body to the outside. The growth rate of the robotic structure is additionally constrained by a “tail” (e.g. a string) attached to the end of its everted material. Control of the internal pressure and string release is required for precise control of the system’s growth. This eversion principle coupled with the compliance of the membranes enables soft growing robots to achieve high extension ratios and navigate through cluttered environments without external surface friction. Specific to our application, it allows for navigation through gaps that are significantly smaller than its width, which is required to insert a sling underneath a patient in a supine position. By using eversion, the flexible membrane can grow into small gaps between the subject’s body and the bed surface without inducing damaging sliding friction to the skin and wounding the patient. The lifting force needed for soft growing robotic structures leveraging eversion to grow into the slanted gap underneath the patient has been studied in past literature [hawkes_soft_2017, blumenschein_modeling_2017, choi_development_2020], and shown to be primarily dependent on the internal air pressure and the angle of the slanted gap.
@@ II-A2 Restricting radial expansion
The ideal geometry for a structure meant to grow underneath patients is one that is wide enough to span the patient’s body and flat enough to avoid discomfort. The inflated geometry of soft everting growing robots, barring external forces, is governed by its wall tension when pressurized. By systematically fabricating the robot body such that it experiences wall tension that restricts its radial expansion into a desired shape, we can design the everting body to have a flatter inflated geometry. The first soft growing robotic structure flat enough to be applicable as a growing sling for patients was described in [choi_development_2020]. This design restricts radial expansion by integrating shafts into the flexible membrane serving as the robot body. The shafts provide an antagonistic moment for the radially expansive forces and are aligned such that they do not interfere with eversion. However, this design results in a loss of all-axis bending flexibility that allows for better conformation of the sling to the body. In our design, we address this by using only flexible fabric. As shown in Fig. 2(a), we bond pairs of points along the robot body’s cross-sectional circumference perpendicular to the direction of its net eversion growth to constrain its radial expansion when inflated. This is akin to how some air mattress designs achieve their flat surface topology. To make this design viable for the eversion principle, the bonded points must allow for a middle channel for the flexible membrane to evert from and sufficiently low capstan friction from the internal materials sliding against each other [blumenschein_modeling_2017, haggerty_characterizing_2019].
@@ II-B1 Fabricating the robot body
To implement restriction of radial expansion, we used thermoplastic-coated fabric (TPU-coated ripstop nylon) due to its exceptional tensile strength and ease of forming robust bonds through heat-sealing the coated sides using an ultrasonic welder (VETRON 5064). The process involved cutting the fabric material that forms the robot body’s perimeter. Subsequently, two strips of fabric were cut to span the net growth length of the robot body, as depicted in Fig. 2(b). Each strip was heat-sealed along two predetermined lines on the perimeter fabric material in the direction of the robot body’s net eversion growth. The distance between these lines was prescribed to constrain the radial expansion by a predetermined amount. The calculation of this distance is detailed in Section III. Finally, the perimeter fabric material was folded to realize the form shown in Fig. 2(b), ensuring that the TPU-coated side contacts itself so it can be heat-sealed along its edges to create an airtight enclosure for the robot body.
@@ II-B2 Creating strong attachments
The fastening points between the Hoyer lift cables and the membrane also creates force concentrations when lifting the human body. While most commercial patient slings have webbing loops sewn onto their edges for the cables to attach to, sewing onto the membrane introduces punctures that lead to air leaks and/or tears. Thus, our prototype implements tubular loops of fabric (same as membrane) welded to the membrane, as shown in Fig. 2(b). The loops are welded inside of the membrane’s outer seam such that the welds are pulled in shear instead of peeling for increased strength. Fabricated as such, each of the loops can bear over 68 kg.
@@ II-B3 Wide base
The base and collar were designed with a long, narrow shape to match the flattened cross-section of the robotic sling, as well as enable the base to be placed under the human’s head. As shown in Fig. 2(b), a hose clamp fastens a set of curved blocks around the collar, which clamp the robotic sling membrane to the collar. Because the normal pressure the hose clamp applies at a given point is proportional to its local curvature, the blocks are designed such that the hose clamp holds a roughly elliptical profile (as opposed to the slot-shaped collar) and applies some normal pressure around the entire perimeter without excessively protruding from the long and narrow shape of the base.
@@ III-B2 Arcs of center channel are coradial
A governing principle for the geometry of inflated structures is that, assuming inextensibility and no flexural rigidity, the membrane is deformed such that the internal volume is maximized ($A_{c}=A_{c,max}$). Thus, $\frac{dA_{c}}{d\theta_{c}}=0$ must be true for any $S_{c}$, $L$, and $\theta_{c}$ (geometric parameters which fully determine the cross-section geometry). If $2\alpha=L$ must be true for $A_{c}=A_{c,max}$ to be true for any $S_{c}$, $L$, and $\theta_{c}$, then $p_{0,y}=p_{ct,y}=p_{cb,y}=0$. The cross-sectional area and its derivative with respect to $\theta_{c}$ are expressed as:
@@ III-C Cross-Section Geometry Model
Additionally, given the fabrication parameters, the full geometry of the cross section can then be determined. The radii of the curved segments of the side and center channels are given as $r_{s}=H_{s}/2$ and $r_{c}=H_{c}/2$, respectively. The centers of the side arcs are $\mathbf{p_{sr}}=\left[p_{sr,x},p_{sr,y}\right]^{T}$ and $\mathbf{p_{sl}}=\left[p_{sl,x},p_{sl,y}\right]^{T}$, where $p_{sr,x}=\frac{w}{2}-r_{s}$, $p_{sl,x}=-\frac{w}{2}+r_{s}$, and $p_{sr,y}=p_{sl,y}=0$. The centers of the top and bottom arcs are already defined previously as $\mathbf{p_{ct}}=\mathbf{p_{cb}}=\mathbf{p_{0}}=\left[0,0\right]^{T}$. The midpoints of the straight segments are given as: $\mathbf{p_{Lr}}=\left[\frac{w_{c}}{2},0\right]^{T}$ and $\mathbf{p_{Ll}}=\left[-\frac{w_{c}}{2},0\right]^{T}$, where the width of the center channel $w_{c}$ is given as
@@ IV-B Effects of Fabrication Parameters on Growth Rate
Recognizing that this design approach limits the space inside the robot body for new material to evert through thus increasing the contact area between the internal everting material and external everted material and resultantly the capstan friction experienced by the system, we performed this experiment to investigate the effects of the fabrication parameters, $S_{c}$ and $L$ on the growth rate of the soft growing robotic structure at a constant pressure. We fabricated three iterations of the robot body using our design approach holding their perimeter distance at a constant $559$ mm but varying the fabrication parameters holding one of each fabrication parameter constant in at least two of the iterations. The iterations are the same as those used in the geometry model validation experiments. Structure 2 holds the same $S_{c}$ as Structure 1 but a smaller $L$ to study the effects of $L$ on growth rate. Structure 3 holds the same $L$ as Structure 1 but a smaller $S_{c}$ to study the effects of $S_{c}$ on growth rate. To provide a basis of comparison to previously presented soft growing robots without added bonds along their cross-sectional perimeter [hawkes_soft_2017], we also fabricated and tested a tubular body structure of the same perimeter distance as the other structures with an unconstrained radial expansion. Each structure was $914$ mm in length and attached to an airtight pressurized base. We use the same pressure regulator and pressure sensor used in the geometry model validation experiments to control and monitor the robot body internal pressure. As a preliminary step, the robot body is everted to its full length, and then inverted by $432$ mm to establish the starting position. Before each trial, the robot body is inverted back to the starting position. For each trial, the robot growth is physically impeded while it is inflated to a starting pressure of 2.07 kPa. Once the starting pressure is reached, the robot growth is unobstructed and the robot is allowed to evert until it stops. To limit the effects of fabric wrinkling on the growth rate, we chose a pressure level sufficiently high enough for the robot body to consistently grow to complete eversion without stopping. We perform three trials for each robotic structure. Each trial was filmed with a high-definition camera at 30 frames per second, and the video was post-processed using MATLAB’s Computer Vision Toolbox to track the growth rate of the robotic structures. The results are presented in Fig. 4(b).
@@ IV-C System Deployment
To demonstrate the effectiveness of our system, the prototype was used to harness and transfer human subjects from a bed to a chair, and then back to the bed. This demonstration was performed with five participants (5 adult males, age 19-32, weight 56-98 kg). As shown in Fig. 5, with the participant lying prone on a bed, a caregiver placed the base with the robotic sling fully inverted underneath their head. The base was inflated to 34 kPa (5 psi) and the robotic sling was everted to grow underneath the participant. Once grown to the full length of their body, the robotic sling was then deflated to assume a sheet-like form, and the caregiver then attached the loops on the robotic sling to the cables on a Hoyer lift. The Hoyer lift was then operated to lift the patient from the bed and move them into a chair, using standard procedures used in caregiving industry. To reverse the transfer, the participant was lifted from the chair and moved back to lay them onto the bed. The robotic sling was then inflated and inverted to remove it from underneath the participant, after which the caregiver removed the base from under the participant’s head. The study was approved by the Massachusetts Institute of Technology’s Internal Review Board, and participants gave informed consent (participants appearing in the supplemental video gave informed consent to allow their faces to be visible). The prototype was successful in completing every step of the demonstration for all participants, in that it automatically placed the robotic sling under the participant’s body without human intervention, it enabled the participant to be lifted and transferred to and from a chair with the same procedure used in caregiving industry, and the robotic sling was then automatically removed from under the participant without human intervention.
@@
As a method of quantifying the comfort of patients while the robot is deployed underneath them, we define an ergonomic index as the inverse of the slope between the highest point of the $A_{s}$ channel and the highest point of the $A_{c}$ channel (denoted in Fig. 3(a)). As intended by the design, a horizontal slot cross-sectional shape allows for a flatter surface for the patient to lay on by minimizing discomfort caused by the rounding of the patient’s back that is present with cylindrical soft growing robotic structures. As defined above, a higher ergonomic index trends towards a concave cross-sectional profile of the robotic sling to better achieve a horizontal slot shape. To investigate the effects the fabrication parameters have on the ergonomic index, we use our geometric model to simulate a sweep of ergonomic indices when holding the structure’s perimeter distance constant while varying the fabrication parameters. The results are presented in Fig. 4(c). The data suggests that, for parameter values that yield physically reasonable shapes, decreasing $L_{c}$ while holding $S_{c}$ constant increases the ergonomic index. It also suggests that decreasing $S_{c}$ while holding $L_{c}$ constant increases the ergonomic index. For each combination of $S_{c}$ and $L_{c}$, there exists a cross-sectional perimeter value that maximizes the ergonomic index which practitioners of this design approach can leverage the analytical model to find.